40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Wed May 19 08:40:47 2021, Stephen, DailyProgress, Cryo vacuum chamber, Damaged workpiece, but otherwise ready to pump down IMG_8743.JPGIMG_8745.JPG
    Reply  Thu May 20 16:40:09 2021, Radhika, DailyProgress, Cryo vacuum chamber, Damaged workpiece, but otherwise ready to pump down IMG_1646.jpegIMG_1651.jpegIMG_1654.jpegIMG_1661.jpeg
Message ID: 2571     Entry time: Thu May 20 16:40:09 2021     In reply to: 2570
Author: Radhika 
Type: DailyProgress 
Category: Cryo vacuum chamber 
Subject: Damaged workpiece, but otherwise ready to pump down 

Following up from Stephen's last post. Today I completed the outstanding tasks he outlined, with the exception of connecting a com cable to record and trend cooldown from the temperature controller. For today's cooldown we are still using the usb flash drive.

The RTD connected to the workpiece (spring clamp) took some wrestling to get stable temp readout, so I had to reclamp it. Other than that, I was able to close up the radiation shields' lids and the chamber lid straight away.

I initially tried pumping down but the backing pump was very loud, not reaching full speed, and pressure wasn't decreasing. Stephen realized the KF joint on the side of the chamber was never sealed up (just wrapped in foil) --> obvious major leak. We didn't have any blanks to seal it up, so I replaced the KF port with a blank flange at the conflat joint.

Ready to pump down once more, I ran into an error message from the turbo pump but performed a power cycle and it disappeared. Pumped down and reached a millitorr before turning on cryocooler.

I plan to pop by tomorrow morning before the cryo meeting and can share consequent updates then.



[Radhika, Stephen]

Good progress toward pumping down, with a setback (impact unknown while we reach out to Karthik).

  • Made final connection of copper braid to baseplate. With the current length, we must pull the folded-over braid completely taught. We are shorting on the aperture of the inner shield, but we appear to have clearance from the vacuum skin (checked by hand at the cryocooler joint, checked by eye at the chamber) and the outer shield. We should be able to resolve these issues with the plan for a long rigid copper bar interfacing directly to the cold head, plus a short copper rope thermal linkage.
  • Completed vacuum flange connections at cryocooler and vacuum gauge on T.
    • See IMG_8743 for overview of current cryocooler connection.
  • Installed Heater on workpiece holder using 2x #4-40 screws, with an indium gasket underneath. Need to check whether we adequately compressed the indium (needs a certain pressure to flow into surface microroughness), but we didn't have any flexibility in the position or bolting arrangement - something to consider for future sample holders.
  • Connected RTDs at Inner Shield (cryo varnish), Outer Shield (cryo varnish), Heater (cryo varnish), and Workpiece (spring clamp). Heater RTD Kapton Tape separating soldered leads needed to be separated, as a short was found at those bare leads. Workpiece RTD connected directly to the workpiece holder since we want to witness the temperature of the sample as directly as possible, and the RTD mounted to the outside of the heater will be overly-sensitive to heat input, yielding a very different time constant from the workpiece (and an inaccurate witness to sample temperature).
    • See IMG_8745 for overview of internal configuration, including RTD positions and thermal linkage to baseplate.
  • Outer and Inner radiation shields were aligned to the copper braid connection to avoid shorting, then to the optical beam path by eye.
  • While making final tightening of workpiece holder, Stephen's allen key slipped, and the installed silicon cantilever was whacked and fractured. The workpiece holder will need to be registered, and the workpiece will need to be swapped out.

The following is the list of remaining actions before we have cooldown data:

  1. Replace workpiece (may require removal of workpiece holder, may require removal and reconnection of RTDs
  2. Check alignment of radiation shield apertures, confirming outer radiation shield is still standing on the 3x G10 spacers
  3. Confirm conductivity of RTDs, both by eye and by temperature controller readout.
  4. Add radiation shield lids, including bolts to inner radiation shield lid.
  5. Bolt down top chamber lid, incremental torquing 1/4-turn at a time with metal-to-metal end result (to achieve full compression of o-ring).
    1. Would be good to wipe down outer surface of o-ring using a wiper and IPA, since there have been a lot of on/off moves since we last pumped down, and we could have picked up particulate introducing a leak path. Check by eye after wipe down.
  6. Pump down system and confirm that we have no leaks (tightening new CF joints would be the first resolution, if there are any issues)
    1. Gauge cables to controller/readout are currently disconnected, need to reconnect.
    2. Can we collect and trend pumpdown data yet from the gauge controller? Might be worth setting this up, if the com cable is already in place as it appears to be.
  7. Turn on cryocooler once vacuum pressure reaches tens of milliTorr or better (should be an hour or two).
    1. Need to start USB datalogging at temperature controller.
    2. RTD names might need updating for consistency with current mounting.
    3. Can we collect and trend pumpdown data yet from the temperature controller? Might be worth setting this up, if the com cable is already in place as it appears to be.

We are both working on adding all of our photos to the photo dump at the ligo.wbridge QIL Cryostat Photo Album. We will then collaborate to add some of the most interesting images to this log!


Attachment 1: IMG_1646.jpeg  1.512 MB  Uploaded Thu May 20 18:28:42 2021  | Hide | Hide all
Attachment 2: IMG_1651.jpeg  1.979 MB  Uploaded Thu May 20 18:29:08 2021  | Hide | Hide all
Attachment 3: IMG_1654.jpeg  2.338 MB  Uploaded Thu May 20 18:29:31 2021  | Hide | Hide all
Attachment 4: IMG_1661.jpeg  1.621 MB  Uploaded Thu May 20 18:29:54 2021  | Hide | Hide all
ELOG V3.1.3-